• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 17:41:55 on May27,2008

    Entry number 239927

    This entry is a followup to: 239885

    keyword=R-HRS Cavity status

    For the cavity monitors on R-HRS, here is the status (see fig 1
    for illustration).

    1. Cav1 is producing strange narrow signals on X,Y,Q. They do look
    different from beam-off, so I suppose it's not a cabling problem.
    Cav2, meanwhile, looks pretty normal (for e.g., see fig 1, top left)

    2. Two channels are saturating. The ADCs work ok up to 8192, then
    jump to 16384. All the channels are close to saturation at 15.8 uA.
    Need to take access and reduce the ADC gate.

    3. Despite these problems there is one combination of striplines
    versus cavities that looks normal (see ellipse, bottom right).
    (Why ellipse ? because of phases, etc). This involves Cav2 and
    the remaining non-saturated channels. So, there is hope.

    Things to do:

    1. Reduce ADC gate. (10 min. access)
    2. Fix Cav1 (issues unknown).
    3. Need a pedestal run (forced gain=0).

    Run plan once these are fixed:

    1. Monitor beam current using HRS trigger. Plan to scan from 10 uA
    down to 50 nA (at 50 nA the BCMs don't work).
    2. At 10 uA, scan X (and later Y) using a steering magnet current I.
    Observe dX/dI.
    3. Repeat #2 with 1 uA, 0.1uA, 50 nA, etc. See if cavities work
    (X,Y,Q track as expected) and measure the resolution.


    FIGURE 1